OpenStudy (anonymous):

What mass of solid is produced by the mixing of aqueous solutions containing 2.00 g AgNO3 (169.9 g/mol) and 4.00 g KBr (119.0 g/mol)?

OpenStudy (preetha):

This sort of problem is called a limiting reagent problem. First write the reaction: AgNO3+KBr-> AgBr(s)+KNO3(aq) 2. Balance it : looks fine 3. Calculate molar masses Bring your periodic table out: Get the accurate molar masses: http://periodic.lanl.gov/index.shtml It is given in this problem. AgNO3 = 107.9 + 15+ 3(16) =171 KBr = 40+80 = 120 4. calculate moles of each moles of AgNO3 = 2g x 1 mol/171 g/mol = 0.011 moles of KBr =4 g x 1 mol/120 g/mol = 0.033 5. Which is the limiting reagent? How many moles of KBr are needed to react with the AgNO3 0.011 moles AgNO3 x 1mol KBr/1mol AgNO3 = 0.011 moles KBr How many moles of AgNO3 are needed to react with 0.033 moles KBr 0.033 moles KBr x 1mol AgNO3/1mol KBr = 0.033 moles AgNO3. 6. But you only have 0.011 moles of AgNO3, so AgNO3 is the limiting reagent and KBr is the excess reagent. 0.011 moles of KBr will react. 7. So moles of AgBr forme 0.011 moles KBr x 1mole AgBr/1mol KBr 8. Mass of AgBr formed Calculate molar mass of AgBr = 108+80 = 188 g/mol mass of solid = 188 g/mol x 0.011 mol = 2.068g

OpenStudy (anonymous):

awesome xplanation.... :))